OpenStudy (anonymous):

evaluate the integral from pi to 0. sin^4 (3t) dt

OpenStudy (amistre64):

sin^4 = sin^2 sin^2 for a start

OpenStudy (amistre64):

sin^2 = 1/2-cos(2t)/2 right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so i did (1/2 (1-cos6t))^2

OpenStudy (amistre64):

id have to do it on paper ...

OpenStudy (anonymous):

jjiji ok

OpenStudy (amistre64):

u = 3t to clear that up .... du = 3 dt; dt = du/3

OpenStudy (amistre64):

i cant recall the steps to well at the moment ... but this is what it amounts to: http://www.wolframalpha.com/input/?i=int%28sin^4%28u%29%2F3%29du

OpenStudy (anonymous):